# A numerical analysis of Araki-Uhlmann relative entropy in Quantum Field Theory

### v2

@@ I Introduction
In this work, we investigate the Araki–Uhlmann relative entropy between a coherent state and the vacuum state of a massive scalar field in (1+1)-dimensional Minkowski spacetime. Utilizing modular theory and the Tomita–Takesaki framework, we develop a numerical approach to explore the positivity of this entropy and its behavior with respect to key parameters, such as mass and the size of spacetime regions.
@@
Additionally, we examine the dependence of entropy on the mass parameter. We demonstrate that, across all cases considered, the entropy decreases as the mass increases. To the best of our knowledge, this finding is both novel and significant, providing new insights into the behavior of entropy in quantum field theory.
@@ IV Numerical setup and results
We start this section by providing some details about the numerical setup we have designed. According to expression (19), the basic integral to be evaluated is $\Delta_{PJ}(f,f^{\prime}_{s}|_{s=0})$. Because deriving a closed analytical expression for the Fourier transform of the test function (24) is challenging, the integral has been directly evaluated in configuration space (see Eq. (34)). The numerical integration has been performed with Mathematica by using the QuasiMontecarlo method with precision set by MaxPoints $=10^{8}$.
@@ IV.1.1 Second alternative: disk-supported test function
Figure 4: We show the bahavior of the Araki-Uhlmann relative entropy as a function of the mass parameter $m$ for $a=2$ and $b=200$ (left panel) and as a function of the radius $a$ for $m=0.1$ and $b=200$ (right panel) for the test function $f_{2}(t,x)$, Eq. (26). For both plots we set $s$ to zero.
@@ V Conclusion
This work presented a detailed numerical investigation of the Araki-Uhlmann relative entropy in QFT, specifically focusing on the relative entropy between a coherent state and the vacuum state for a free massive scalar field in $1+1$ Minkowski spacetime. The study employed a numerical setup based on Tomita-Takesaki modular theory and explicit test functions profiles, and has revealed a universal behavior governing the Araki-Uhlmann relative entropy dependence on key parameters.
Across all test function profiles studied — including those supported in the right Rindler wedge, a disk, a smooth diamond, and a vertical strip — the relative entropy consistently decreases as the mass parameter $m$ increases. This trend aligns with physical intuition, as larger masses suppress quantum fluctuations, leading to a reduction in entanglement contributions. Conversely, the relative entropy exhibits a monotonic increase as the spatial extent of the test function’s support grows, whether characterized by the slicing parameter $\alpha$, the radius $a$, or analogous measures in other cases. This behavior confirms the expected result that larger regions capture more entanglement, reinforcing the positivity of the relative entropy and the interpretation of it as a robust measure of quantum distinguishability and entanglement in QFT.
Importantly, while the specific numerical values depend on the precise choice of the test function, the overarching trends remain unchanged, demonstrating the universality of these entropy properties.

### v4

@@ I Introduction
In this work, we investigate the Araki–Uhlmann relative entropy between a coherent state and the vacuum state of a massive scalar field in (1+1)-dimensional Minkowski spacetime. Despite the conceptual significance of the relative entropy, explicit evaluations in field-theoretic settings often demand careful numerical treatment. To this end, by using modular theory and the Tomita–Takesaki framework, we develop a numerical approach to explore the positivity of this entropy and its behavior with respect to key parameters, such as mass and the size of spacetime regions.
@@
Additionally, we examine the dependence of entropy on the mass parameter. We demonstrate that, across all cases considered, the entropy decreases as the mass increases. To the best of our knowledge, this finding is both novel and significant, providing new insights into the behavior of entropy in QFT.
The principal findings can be summarized as follows:
  • We have shown in an explicit way that the behavior of the relative entropy does not depend on the specific choice of the test functions. This is done by showing, through a numerical approach, that the behavior of the entropy remains unchanged upon using four different types of test functions.
  • Relative entropy decreases monotonically with increasing mass. While this trend aligns with theoretical expectations,  we are not aware of any existing formal proof of this result. Also here, the same decreasing behavior with respect to the  mass parameter  is common to all employed  test functions. Our results thus provide the first explicit confirmation across different test functions.
  • Finally, we provide an explicit check of the monotonic increase of the entropy with the size of the spacetime regions under considerations.
These findings reinforce our current understanding of the Araki-Uhlmann entropy and offer concrete numerical evidence supporting its robustness in free field theories.
@@ IV Numerical setup and results
We start this section by providing some details about the numerical setup we have designed. According to expression (19), the basic integral to be evaluated is $\Delta_{PJ}(f,f^{\prime}_{s}|_{s=0})$. Because deriving a closed analytical expression for the Fourier transform of the test function (24) is challenging, the integral has been directly evaluated in configuration space (see Eq. (35)). The numerical integration has been performed with Mathematica by using the Quasi-Montecarlo method with precision set by MaxPoints $=10^{8}$.
More specifically, the numerical computation of the Araki-Uhlmann relative entropy requires the evaluation of a spacetime integral of the form
${\cal J}(m,\alpha)=\int d^{2}x\;d^{2}{x^{\prime}}\;f(t,x)\Delta_{P}J(t-t^{\prime},x-x^{\prime})f^{\prime}_{s}(t^{\prime},x^{\prime})\Big{|}_{s=0}\;,$ (25)
where, for definiteness, we adopt the test function specified in Eq. (24). Expression (25) represents a four-dimensional integral, with the domain of integration determined by the support of the test function $f(t,x)$. To generate a plot analogous to the left panel of Fig. (2), one begins by selecting a fixed value for the parameter $\alpha$, for instance $\alpha=6$. The integral is then numerically evaluated using the Quasi-Monte Carlo method implemented in Mathematica, over a set of points uniformly distributed in the interval [0,5.6], as illustrated in Fig. (2). The mass parameter $m$ is discretized with a step size of 0.01. This procedure yields an initial discrete (dotted) dataset. A continuous plot is subsequently obtained by interpolating the resulting data points. This summarizes the essential numerical strategy adopted in the computation.
@@ IV.1.1 Second alternative: disk-supported test function
Figure 4: We show the behavior of the Araki-Uhlmann relative entropy as a function of the mass parameter $m$ for $a=2$ and $b=200$ (left panel) and as a function of the radius $a$ for $m=0.1$ and $b=200$ (right panel) for the test function $f_{2}(t,x)$, Eq. (27). For both plots we set $s$ to zero.
@@ V Conclusion
This work presented a detailed numerical investigation of the Araki-Uhlmann relative entropy in quantum field theory, focusing on the case of a coherent state relative to the vacuum in a free massive scalar field in $1+1$-dimensional Minkowski spacetime. The analysis employed a framework grounded in Tomita–Takesaki modular theory and systematically explored various classes of test functions with compact support, including configurations localized in the right Rindler wedge, a disk, a smooth diamond, and a vertical strip.
Across all test function profiles, the relative entropy was found to exhibit two robust behaviors: a monotonic decrease with the mass parameter $m$, and a monotonic increase with the size of the spacetime region defined by the test function’s support (as quantified by parameters such as the slicing parameter $\alpha$, the disk radius $a$, or their analogues in other geometries). These trends are consistent with physical intuition—larger mass suppresses quantum fluctuations, reducing entanglement, while larger regions capture more quantum correlations.
Importantly, although these behaviors are theoretically anticipated, to our knowledge no formal proof exists in the literature. Thus, this work provides an explicit numerical exploration and confirmation of these expectations within the framework of free field theory. The observed universality—i.e., the independence of qualitative features from the specific choice of test function—underscores the structural robustness of the Araki-Uhlmann relative entropy as a measure of quantum distinguishability and entanglement.
@@
Finally, we briefly address the highly nontrivial problem of evaluating the Araki-Uhlmann relative entropy in the context of interacting QFTs. In principle, for theories defined on $1+1-$dimensional Minkowski spacetime, one may exploit the powerful framework of bosonization, particularly in models such as the massive Thirring model, which is dual to the sine-Gordon theory [45, 46]. The availability of exact correlation functions in such settings offers a promising avenue to analyze the dependence of the relative entropy on the interaction coupling constant. Any significant advancements in this direction will be communicated in due course.